A father is 6 times as old as his son. Three years ago , the father was 9 times as old as his son. Find the present age of each.
Answer by amarjeeth123(569)

Let the son's age be x.
The father's age is 6x.
Three years ago the father's age is (6x-3)
Three years ago the son's age is x-3.
6x-3=9(x-3)
6x-3=9x-27
3x=24
x=8
The father's age is 48 years and the son's age is 8 years.